OpenStudy (luigi0210):

Find the derivative:

OpenStudy (luigi0210):

\[y=(\theta^2+\sec \theta+1)^3\]

OpenStudy (abb0t):



terenzreignz (terenzreignz):

Chain rule comes to mind...

OpenStudy (abb0t):

\(\frac{d}{dx} [a]^n = \frac{d}{dx}n[a]^{n-1} \times \frac{d}{dx}(a)\)

OpenStudy (abb0t):

or mayb you should think of \(a\) as \(f(x)\) instead of single term \(a\) it kind of looks like power rule

terenzreignz (terenzreignz):

Let's let \[\Large u = \theta ^2 +\sec\theta + 1\] So that we have \[\Large y = u^3\] Then \[\Large \frac{dy}{d\theta}=\frac{dy}{du}\cdot \frac{du}{d\theta}\] as per the chain rule...

OpenStudy (abb0t):

\(f(\theta)\)**

terenzreignz (terenzreignz):

Think of this \[\Large y= (\color{blue}{\theta^2 + \sec\theta+1} )^3\]as a single entity first, and differentiate, using the power rule: \[\Large \frac{d}{dx}x^n = nx^{n-1}\]

OpenStudy (luigi0210):

so first part: \[y'=3(\theta^2+\sec \theta+1)^2*2 \theta+secxtanx\]

terenzreignz (terenzreignz):

Better enclose these in parentheses instead...and it's \(\large \sec\color{red}\theta \tan \color{red}\theta\) not x lol... \[\Large y' = 3(\theta^2 + \sec\theta + 1)^2(2\theta+ \sec\theta\tan\theta )\] Yes, parentheses matter :3

OpenStudy (psymon):

Just don't drop the parenthesis around (2theta + sectan) and I was beat to it, haha.

OpenStudy (luigi0210):

Oops, right.. thanks guys :)

OpenStudy (abb0t):

\(\huge \theta\)